OpenStudy (anonymous):

Tickets for a school event were $1.75 for students and $3.25 for parents. If a total of 72 tickets were sold for $195.00, how many student and parent tickets were sold?

OpenStudy (campbell_st):

let s = students and p = parents number of people s + p = 72 (1) income from attendees is 1.75s + 3.25p = 195 (2) from equation (1) make s the subject and then substitute into equation (2) and solve for p.

OpenStudy (anonymous):

wait i am so confused what to do from here? @campbell_st

OpenStudy (anonymous):

@campbell_st ???

OpenStudy (campbell_st):

ok.... let s = 72 - p.... thats from the 1st equation does that make sense